Think of all of the purchases you have made over the last week:

  1. What percentage of the purchases was based on utilitarian consumption motives?
  2. What percent was based on hedonic consumption motives? Describe your most recent hedonic consumption experience.
Answer #1

Ans) Utilitarian consumption is consumption on things which are necessary for survival. Like food, water, clothing, shelter, medical care etc.

After consumption of utilitarian goods and services if you are left with funds, you spend it on leisure goods and items. This is called hedonic consumption. Example- music, vacation, shopping for pleasure etc.

Recently I bought food and medicine where there were 20 food items and 3 medicines. Also I bought a Video game and a earphone.

So total goods that I bought is 25 out of which 23 is utilitarian consumption (food and medicine) and 2 is hedonic consumption (video game and earphones).

Therefore percentage of purchase of utilitarian consumption = (23÷ 25) × 100 = 92%

And percentage of purchase of hedonic consumption = (2÷25)×100 = 8%

My experience with my headphone and video game was great. The earphones have very good sound quality and bass, they are very comfortable and their bluetooth range is 15 meters which gives me additional comfort that I do not have to hold my phone all the time .

My video game quality is also very nice, the game is different and I enjoy playing it.